OpenStudy (anonymous):

solve each equation. 3x-17=22 (I got x=13. Not sure if that was right. I added 17 on each side and that gave me 3x=39 and so I divided it by 3. and got x= 13. ) 13(2x)+20=6(5x+2) (so on this one I know you have to distribute the 6 so it would make that 30x+12,,Right? However looking at the other side would I distribute that too?) 8 ( x - 5 ) = 2x-6 --- 2 (On this one,, I don't even know how to go about it,,,)

OpenStudy (anonymous):

For your first equation, you can check your answer by plugging 13 back in for x. 39 - 17 does equal 22 so you are correct.

OpenStudy (anonymous):

13(2x)+20=6(5x+2) 26x + 20 = 30x + 12 -4x = -8 x = 2

OpenStudy (anonymous):

For your last equation, multiply both sides by two to remove the division and then solve the same as the second one.

OpenStudy (anonymous):

thank you so much!!